Qhat are the measures of two complementary angles if the difference in the measures of the two angles is 12?
Answer by Earlsdon(6294)

Recall that if two angles are complementary, their sum is 90 degrees.
If you let the two angles be x and y, then:
x + y = 90.
In the problem description, the difference of these two angles is 12 degrees.
So you can write:
x - y = 12 By adding y to both sides, you can rewrite this as:
x = y+12 Now substitute this for x in the first equation.
(y+12) + y = 90 Simplify and solve for y.
2y+12 = 90 Subtract 12 from both sides.
2y = 78 Finally, divide both sides by 2.
y = 39 degrees, and:
x = y+12
x = 39+12
x = 51 degrees.
x+y = 90
39 + 51 = 90
90 = 90
x-y = 12
51-39 = 12
12 = 12